# An abstract framework for heterogeneous coupling: stability, approximation and applications

### v1

@@ 4. Solving the resulting discrete problem
$C(V_{h},d,\sigma)=\|\Pi_{d}\|_{V\to V}\left(1+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}\|_{V\to V}\right).$
4.1. Strong form of the iterative solver
Let us give some detail on how the different steps involved in the reduction of the discrete problem and in the application of the preconditioner can be split as composition of a small numbers of linear operators, which are the ones that need to be implemented. To this aim we introduce the Riesz operators relative to the two scalar products $\sigma$ and $d$, namely $\Sigma:\Lambda_{\delta}\to\Lambda_{\delta}^{\prime}$ and $D:V_{h}\to V_{h}^{\prime}$ be defined as
$\langle\Sigma\lambda,\mu\rangle=\sigma(\lambda,\mu),\ \forall\lambda,\mu\in\Lambda_{\delta},\qquad\langle Dv,w\rangle=d(v,w),\ \forall v,w\in V_{h}.$
Moreover we let $G_{\delta}:Z\to\Lambda_{\delta}^{\prime}$ denote the restriction to $Z$ of the operator $B_{\delta}$:
$\langle G_{\delta}z,\lambda\rangle=\langle B_{\delta}z,\lambda\rangle=b(z,\lambda),\ \forall z\in Z,\ \lambda\in\Lambda_{\delta}.$
Remark that we need to distinguish between the operators $B_{\delta}$ and its restriction $G_{\delta}$ as we will need to take the adjoint operators. Some standard linear algebra manipulations allow to see that the projectors $\Pi_{d}$ and $\Pi_{\sigma}$ verify respectively
$\Pi_{d}=D^{-1}B_{\delta}^{T}(B_{\delta}D^{-1}B_{\delta}^{T})^{-1}B_{\delta},\qquad\Pi_{\sigma}=1-\Sigma^{-1}G_{\delta}(G_{\delta}^{T}\Sigma^{-1}G_{\delta})^{-1}G_{\delta}^{T}.$
while the reduced equation (4.2) in strong form reads
(4.6) $\mathcal{I}^{T}S_{h}\mathcal{I}\eta=\mathcal{I}^{T}g-\mathcal{I}^{T}S_{h}\lambda^{0},$
@@ 4.1. Strong form of the iterative solver
$\lambda^{0}=-\Sigma^{-1}G(G^{T}\Sigma^{-1}G)^{-1}f.$
The pseudo inverse $B_{\delta}$ can be realized as
$B_{\delta}^{+}=D^{-1}B_{\delta}^{T}(B_{\delta}D^{-1}B_{\delta}^{T})^{-1},$
and the linear operator corresponding to the bilinear form $m$ is
$M=(B_{\delta}^{+})^{T}A_{\delta}B_{\delta}^{+}.$
Finally, exploiting Proposition 4.1 we can write
$\widehat{M}=\Pi_{\sigma}(B_{\delta}^{+})^{T}A_{\delta}B_{\delta}^{+}\Pi_{\sigma}^{T}.$
Remark that, while the operators $A_{\delta}$, $B_{\delta}$ and $G_{\delta}$ are substantially given by the continuous problem, there is, a priori, a certain freedom in the choice of the operators $D$ and $\Sigma$ (or, equivalently, of the scalar products $d$ for $V_{h}$ and $\sigma$ for $\Lambda_{\delta}$). These have to be chosen in such a way that the resulting preconditioner $\widehat{M}$ is cheap (or easily parallelizable), and that, at the same time, the norms of $\Pi_{d}$ and $B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}$ stay bounded. How this can be done depends on the actual problem considered. Remark that choosing $\sigma$ and $d$ as the natural scalar products for $\Lambda$ and $V$ would immediately result in the boundedness of the above mentioned operators but would possibly lead to expensive preconditioner.
5. Applications
The above framework has multiple possible applications. The most obvious one is of course to design domain decomposition solvers for large scale problems. However it can also be used for the design of multiphysics, multiscale or mixed-dimensional problems. Here we show some typical model situations that enter the framework. We develop the domain decomposition case in the general setting of a multi-domain coupling. In the other examples, for the sake of notational simplicity, we consider model situations where only two or three domains are coupled. The extension to the multi-domain case is straightforward using the arguments of the first example.
5.1. Domain decomposition with Neumann or Dirichlet coupling
Let $\Omega\subset\mathbb{R}^{2}$ be a polygonal domain with boundary $\Gamma$. We consider the following simple model problem: given $f\in L^{2}(\Omega)$, find $u$ satisfying

### v2

@@ 5.2. Domain decomposition with Dirichlet coupling
(5.16) $a^{k}(u^{k},\phi^{k};v^{k},\psi^{k})=\int_{\Omega^{k}}\kappa\nabla u^{k}\cdot\nabla v^{k}-\int_{\Gamma^{k}}\phi^{k}v^{k}\\ +\int_{\Gamma^{k}}u^{k}\psi^{k}+[\mathcal{D}^{k}u^{k}-\gamma_{\Gamma^{k}}^{*}\phi^{k},\mathcal{D}^{k}v^{k}-\gamma_{\Gamma^{k}}^{*}\psi^{k}]_{-1,\Omega^{k}},$
and with $\bar{\kappa}^{k}=\fint_{\Omega^{k}}\kappa$,
$b^{k}(\phi^{k};\mu)=\bar{\kappa}^{k}\int_{\Gamma^{k}}\phi^{k}\mu$
$\langle F;v^{k},\psi^{k}\rangle=\int_{\Omega^{k}}fv^{k}+[f,\mathcal{D}^{k}v^{k}-\gamma_{\Gamma^{k}}^{*}\psi^{k}]_{-1,\Omega^{k}}+\int_{\Gamma^{k}}v^{k}\fint_{\Gamma^{k}}\zeta^{k}$
$\langle G,\mu\rangle=-\sum_{k}\int_{\Gamma^{k}}\mu\fint_{\Gamma^{k}}\zeta^{k}.$
In the above expression, $\mathcal{D}^{k}:H^{1}(\Omega^{k})\to H^{1}(\Omega^{k})^{\prime}$ is the linear operator corresponding to the semi-scalar product induced by $a^{k}$,
$a^{k}(u,v)=\langle\mathcal{D}^{k}u,v\rangle,$
$\gamma_{\Gamma^{k}}:H^{1}(\Omega^{k})\to H^{1/2}(\partial\Omega^{k})$ is the trace operator, $\gamma_{\Gamma^{k}}^{*}$ denoting its adjoint, and $\cdot,\cdot]_{-1,\Omega^{k}}$ stands for the $H^{1}_{\text{av}}(\Omega^{k})^{\prime}$ scalar product. We recall that the bilinear form $a^{k}$ does not need to play any role in the design of the black box local solvers (see Remark 3.13). It will however play a role in the design of the preconditioner. To see that the formulation (5.14)–(5.15) is indeed consistent with (5.12)–(5.13) consider the Riesz map $\Psi:(H^{1}(\Omega^{k}))^{\prime}\to H^{1}(\Omega^{k})$ and define
$v,w]_{-1,\Omega^{k}}=\langle v,\Psi w\rangle,\quad\mbox{ for }v.w\in(H^{1}(\Omega^{k}))^{\prime}.$
It then follows that if we set $w^{k}=\mathcal{D}^{k}v^{k}-\gamma_{\Gamma^{k}}^{*}\psi^{k}$ and $\tilde{w}^{k}=\Psi w^{k}$,
$\mathcal{D}^{k}u^{k}-\gamma_{\Gamma^{k}}^{*}\phi^{k},w^{k}]_{-1,\Omega^{k}}=\left<\mathcal{D}^{k}u^{k}-\gamma_{\Gamma^{k}}^{*}\phi^{k},\tilde{w}^{k}\right>=a^{k}(u^{k},\tilde{w}^{k})-\left<\phi^{k},\tilde{w}\right$>.
Using the definition of $\phi^{k}$ we have
$a^{k}(u^{k},\tilde{w}^{k})-\left<\phi^{k},\tilde{w}\right>=a^{k}(u^{k},\tilde{w}^{k})-\left<\zeta^{k},\tilde{w}^{k}\right>+\int_{\Gamma^{k}}\tilde{w}^{k}\fint_{\Gamma^{k}}\zeta^{k}\\ =\int_{\Omega^{k}}f\tilde{w}^{k}+\int_{\Gamma^{k}}\tilde{w}^{k}\fint_{\Gamma^{k}}\zeta^{k}=[f,\mathcal{D}^{k}v^{k}-\gamma_{\Gamma^{k}}^{*}\psi^{k}]_{-1,\Omega^{k}}+\int_{\Gamma^{k}}v^{k}\fint_{\Gamma^{k}}\zeta^{k}.$
It is not difficult to check that the modified system has a trivial kernel. Indeed, for $F=0$, testing (5.14) with $(v^{k},\psi^{k})=(u^{k},\phi^{k})$ we obtain that
$\int_{\Omega^{k}}|\nabla u^{k}|^{2}+\|\mathcal{D}u^{k}-\gamma^{*}_{\Gamma^{k}}\lambda^{k}\|^{2}_{-1,\Omega^{k}}=0$
which implies that $u^{k}$ is a constant (and, consequently, that $\mathcal{D}u^{k}=0$) and $\gamma_{\Gamma^{k}}^{*}\lambda=0$. This last equation implies that $\lambda=0$. Then, testing (5.14) with $(v^{k},\psi^{k})=(0,1)$ we obtain
$\int_{\Gamma^{k}}u^{k}=0,$
which finally implies $u^{k}=0$.
We now consider a discretization of $\Lambda=H^{1}_{0}(\Omega)|_{\Sigma}$ by conforming degree $l$ finite elements. As far as the local solvers are concerned, we point out that we essentially require them to provide us with approximate Dirichlet to Neumann mappings. Assumption 3.3 reduces to asking that such Dirichlet to Neumann mappings are sufficiently precise. Let us then focus on constructing the preconditioner. As in our formulation the control of $\lambda$ is obtained through the test function $\psi$, we can take $V_{\delta}^{k}$ of the form $V_{\delta}^{k}=\{0\}\times\Phi^{k}_{\delta}$, with $\Phi^{k}_{\delta}$ such that
$\inf_{\lambda\in\Lambda}\sup_{\phi\in\Phi^{k}_{\delta}}\frac{\int_{\Gamma^{k}}\phi\lambda}{|\lambda|_{1/2,\Gamma^{k}}\|\phi\|_{-1/2,\Gamma^{k}}}\gtrsim 1,\qquad\text{with}\quad\|\phi\|_{-1/2,\Gamma}=\sup_{w\in H^{1}(\Omega^{k})}\frac{\int_{\Gamma^{k}}\phi w}{\|w\|_{1,\Omega^{k}}}.$
We know that, provided the mesh is quasi uniform, such inf sup conditions are satisfied for $\Phi_{\delta}^{k}=\Lambda|_{\Gamma^{k}}$. We can then take any $\Phi^{k}_{\delta}$ with $\Phi_{\delta}^{k}\supseteq\Lambda|_{\Gamma^{k}}$. It will be convenient to take $\Phi^{k}_{\delta}$ as the space of degree $l$ finite elements, discontinuous at the subdomain vertices (in 2D) or at the wirebasket (in 3D), and continuous everywhere else. As $\ker A^{k}$ is trivial, we have that $\widehat{\Lambda}_{\delta}=\Lambda_{\delta}$ and we can set $\Pi_{\sigma}=\mathbf{1}_{{\Lambda_{\delta}}}$. We can define the bilinear form $\widetilde{a}$ as
$\widetilde{a}^{k}(\phi,\psi)=\sum_{F\text{ face of }\Omega^{k}}\left(a^{k}_{F}(\phi,\psi)+H^{(d-1)}\left(\fint_{F}\phi\right)\left(\fint_{F}\psi\right)\right)$
with, for all face $F$ of the subdomain $\Omega^{k}$,
$a^{k}_{\color[rgb]{0,0.88,0}F}(\phi,\phi)\simeq\sup_{{\zeta\in H^{1/2}(F)}\atop{\int_{F}\zeta=0}}\frac{\int_{F}\phi\zeta}{|\zeta|_{1/2,F}}$
The following Lemma can be obtained by a dual argument to the one underlying the proof of Lemma 5.3.
Lemma 5.5.
For all $\phi\in\Phi^{k}_{\delta}$ it holds that
$a(0,\phi;0,\phi)\lesssim{\widetilde{a}(\phi,\phi)}\lesssim\log(H/\delta)^{2}a(0,\phi;0,\phi).$
In this case, Assumption 4.7 is not satisfied, so we will have to rely on Theorem 4.4 for estimating the condition number. The bilinear form $\widetilde{a}$ is, by construction, split as a coarse component plus a block diagonal component, so that it can be inverted efficiently in parallel. We then let $d=\widetilde{a}$, which automatically yields $c^{\flat}\simeq C^{\flat}\simeq 1$, while Lemma 5.5 yields $\kappa_{*}=1$ and $K^{*}=\log(\delta/H)^{2}$. Using the choices of $\widetilde{a}$, $d$ and $V_{\delta}^{k}$ defined above we can construct $\widehat{M}^{-1}$ following section 4.1. The resulting preconditioner, which is, to our knowledge, novel, is similar to the substructuring preconditioner by Bramble, Pasciak and Schatz, but with a different treatment of the wirebasket. Theorem 4.4 yields the bound
$\kappa(\widehat{M}^{-1}S_{h})\lesssim|\log(H/\delta)|^{2}$
Remark 5.6.
In a finite element context the former domain decomposition formulation (5.2)–(5.3) was first discussed in [38] and typically results in mortar methods [5, 4, 45], or alternatively FETI methods, see [42] and references therein. Remark that, in the framework of the latter domain decomposition formulation (5.12)–(5.13), Problem 3.1 substantially coincides with the three fields domain decomposition method [19] as well as the natural domain decomposition method with non matching grids proposed in [39]. A coupling using the Dirichlet trace as hybrid variable can also be introduced for mixed finite element method, [1]. On the discrete level any method for the weak imposition of Dirichlet conditions may be used to realize (5.12)–(5.13). In particular, one may eliminate the local multiplier by applying Nitsche’s method in the spirit of [40], resulting in the hybridised method proposed in [26]. Note however, that if the multiplier is eliminated, in general the coupling scheme can not be used in a black box fashion, since information on both the Dirichlet and Neumann traces are required. Both these methods have been applied in the context of multi scale methods, see [32] and [37], with analysis of a preconditioner in the second reference.
Acknowledgement
EB was partially funded by the EPSRC project EP/W007460/1, Integrated Simulation at the Exascale: coupling, synthesis and performance. SB is member of the Gruppo Nazionale di Calcolo Scientifico – Istituto Nazionale di Alta Matematica (GNCS–INdAM) and was partially funded by D34Health - Digital Driven Diagnostics, prognostics and therapeutics for sustainable Health care initiative (PNC0000001) and by MUR PRIN 2022 PNRR, grant P2022BH5CB (funded by NextGeneration EU).
Appendix A Proof of Theorem 4.4
$\frac{1}{c_{*}}\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{m^{*}({\widetilde{\varphi}},{\widetilde{\varphi}})}}\lesssim C^{*}\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}$,
@@ Acknowledgement
$c_{*}=\sqrt{K^{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}\left(1+\|B_{\delta}^{+}Q_{\sigma}^{*}B_{\delta}\|_{\widetilde{a}\to\widetilde{a}}\right),\qquad C^{*}=\frac{1}{\sqrt{\kappa_{*}}}.$
We start by observing that
(A.1) $\kappa_{*}\|v\|_{V}^{2}\lesssim\|v\|_{\widetilde{a}}^{2}\lesssim K^{*}\|v\|_{V}^{2}.$
Moreover, we have that
$\|v\|_{\widetilde{a}}\lesssim|v|_{\widetilde{a}}\qquad\text{ for all $v\in\ker B_{\delta}$}.$
We start by observing that, letting ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$, we can split $B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}}\in V_{\delta}$ ($\mathcal{K}$ defined by (4.4)) as $B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}}=z_{\widetilde{\varphi}}+w^{\perp}_{\widetilde{\varphi}}$, with $z_{\widetilde{\varphi}}\in Z$, and $w_{\widetilde{\varphi}}^{\perp}\in V_{\delta}$ $V$-orthogonal to $Z$. As $z\in\ker A=\ker A^{T}$, we have